# Impact of boiling liquid droplets: Vapor entrapment suppression

### v1

@@ Abstract
In this work we replace gas by vapor, by focusing on the impact of boiling liquid droplets, that is, droplets that are at their boiling point, in thermal equilibrium with the vapor that surrounds them. More specifically, we want to focus on the entrapment of vapor inside the dimple formed below the droplet due to pressure accumulation. A related subject is that of droplets impacting in (or close to) the well known Leidenfrost regime [14], where due to evaporation a droplet can levitate over a heated solid upon impact. These have been performed in atmospheric [15, 16, 17, 18] and reduced pressure conditions [19]. Unlike this previous work, we do not use large temperature differences to trigger rapid phase change, since the liquid will be in thermodynamical equilibrium with its own vapor, even before impact. When in such a state, minute energy exchanges of the droplet with its surroundings can trigger phase change [20]. Moreover, unlike the Leidenfrost impact work where phase change is limited to the evaporation of the liquid due to high substrate temperatures, in this boiling liquid state the condensation of vapor due to pressure locally rising above the vapor pressure of the liquid is also possible. We will show that the latter may lead to dramatic effects.
@@
In order to generate droplets in equilibrium with their own vapor and record their impact onto a substrate, we have developed a highly versatile setup, using HFE-7000 as a working fluid, which has an appreciable vapor pressure ($p_{V}=0.57$ bar) at room temperature. Here we provide a general description of the setup (see Fig. 1), while a more detailed description can be found in the Supporting Material.
@@
First, a saturated liquid-vapor environment at temperature $T_{0}\approx 20$ C is created by allowing HFE-7000 to evaporate inside a hermetically sealed, vacated, and temperature-controlled chamber. Enough liquid is allowed into a reservoir (Res1) to ensure the chamber remains saturated with vapor during the entire experiment. Additional HFE-7000 liquid is allowed into a smaller reservoir (Res2), located higher in the setup, connected to a needle (N) with its tip slightly below the liquid level, allowing droplets to fall down and impact a sapphire prism (P). The height between needle and substrate can be varied to obtain an impact velocity $U_{0}$ between $0.2$ m/s and $1.5$ m/s. The substrate temperature $T_{s}$ is controlled between $T_{0}$ and $40$ C.
Side views of the impact are recorded at 16,000 fps and are used to measure the impact speed $U_{0}$ and initial radius $R_{0}$ of the droplets. Bottom views are captured at frame rates between $96,000$ and $200,000$ fps by illuminating the top surface of the prism internally with a collimated monochromatic light source (CLS2) from below, such that it totally reflects. If a droplet comes within the evanescent wave distance from the surface, then light is partly, and upon contact fully transmitted into the droplet, which results in dark regions on the otherwise bright surface. This technique is known as frustrated total internal reflection (TIR) and has been extensively used in the past [23, 24, 25, 17].
@@
We have performed an extensive experimental campaign, divided in two parts. For the first part we have performed experiments where boiling droplets, falling at different velocities, impact a slightly heated surface. The latter is done to avoid the formation of a liquid film on top of the substrate that would interfere with the impact. For simplicity, we will refer to the experiments in this first part as isothermal experiments.
@@
Figure 2: (A) TIR image sequence (after background subtraction) of a series of HFE-7000 droplets with radius $R_{0}\approx 0.6$ $m$ in a vapor atmosphere at $T_{0}\approx 22$ C; the droplets impact a sapphire substrate heated to a temperature difference $\Delta T<2$ C with impact velocities $U_{0}=0.24$, $0.38$, $0.51$, $0.74$ m/s. At low impact velocity a vapor bubble is entrapped, but at high impact velocities this entrapment is suppressed. (B) TIR image sequence (after background subtraction) of a series of droplets with radius $R_{0}\approx 0.6$ $m$ in a vapor atmosphere at $T_{0}\approx 22$ C impacting with velocity $U_{0}\approx 0.7$ m/s onto a sapphire substrate heated to an increasing temperature difference $\Delta T=2.4$, $6.4$, $11.3$, $17.6$ K (with respect to $T_{0}$) At this impact velocity, the entrapment of vapor is suppressed for small $\Delta T$, but the entrapment can be recovered by heating the substrate to a higher temperature. In all cases, time is measured with respect to the first TIR frame in which the droplet becomes visible.
@@
Figure 3: (A) Azimuthally averaged inner ($r_{b}$, blue points) and outer ($r_{d}$, orange points) radius evolution for an experiment at $U_{0}=0.38$ m/s, with images available in the Supporting Material, Fig. S1B. The orange line is obtained from a linear fit to $r_{d}^{2}$ plotted against $t$ and the blue line is a linear fit for $r_{b}$. The intersection of the two curves provides the initial radius $r_{0}$ and also the initial impact time. The entrapment time $t^{*}$ is then measured from the inner radius fit as the time it takes for $r_{b}(t)$ to go from $r_{0}$ to 0. (B) Initial entrapment radius $r_{0}$ as a function of the impact velocity $U_{0}$ for a set of experiments with a small temperature difference of the impact surface with respect to the ambient temperature ($\Delta T=1.77\pm 0.49$ K). The points represent the mean of a minimum of three repetitions and the bars display the total error. (C) entrapment time $t^{*}$ as a function of the impact velocity $U_{0}$, for the same set of experiments.
@@
In the next few paragraphs, through a simplified model, we find a quantitative explanation for the observed suppression of vapor bubble entrapment for increasing impact velocity. We begin considering an impacting droplet surrounded by a non-condensable gas. The pressure difference under such droplet is described by the lubrication flow of the gas as the droplet nears the surface [10]. As argued in that work, we can estimate the pressure build up below the droplet from the axisymmetric Stokes equation in the lubrication layer as
@@
where $\mu_{g}$ is the dynamic viscosity of the gas, $h$ is the thickness and $\ell\sim\sqrt{R_{0}h}$ the radial extent of the lubrication layer, where the latter is estimated geometrically from the approach of a spherical droplet to a planar substrate, and the gas velocity $U_{g}$ is related to the impact velocity by continuity, $U_{g}h\sim U_{0}\ell$. By equating this $\Delta P$ with the pressure needed to decelerate the liquid, namely $\rho_{L}(\partial\phi/\partial t)\sim\rho_{L}U_{0}^{2}\ell/h$, one obtains a scaling relation for the dimple height $h_{d}$ in terms of the Stokes number $\mathit{St}\equiv\rho_{L}U_{0}R_{0}/\mu_{g}$
@@
For our case, however, the surrounding phase is vapor and this means that an increase in pressure can trigger condensation: As soon as the pressure in the vapor layer rises, the vapor will try to condense to return to the equilibrium state. For simplicity, let’s now consider a one-dimensional system composed of a solid substrate, an incoming liquid surface and an intermediate vapor layer. During pressure build up, vapor will try to condense and produce latent heat. Since simultaneously the temperature in the vapor rises, a thermal boundary layer will develop in the liquid, which will allow for transport of the latent heat into the liquid. Now, writing down the mass-energy balance at the liquid-vapor interface [29]
@@
where $R_{s}$ is the specific gas constant of the vapor phase and the equilibrium state variables of the vapor are denoted with subscript 0: $P_{V,0}$, $\rho_{V,0}$, and $T_{0}=T_{V,0}$. Finally, the boundary layer thickness $\delta_{th}$ can be approximated as $\delta_{th}\sim\sqrt{\pi\alpha_{L}\Delta t}$, with $\alpha_{L}$ the thermal diffusivity of the liquid and the time scale $\Delta t\approx\epsilon\frac{R_{0}}{U_{0}}$, where $\epsilon$ should be approximated from evaluating the time that the pressure rise in the lubrication layer starts being appreciable. We take $\epsilon\approx 0.02$, corresponding to $h\approx 0.02R_{0}$, which is consistent with BI simulations (see next Section). Substituting everything into the expression for the condensation speed, and dividing by $U_{0}$ we obtain
@@
Now, the threshold condition $U_{cond}/U_{0}=1$ can be computed by turning the approximate result (8) into an equation using a multiplicative numerical constant $K_{0}$ in the right hand side and subsequently inserting literature values for the transport coefficients for HFE-7000 at $T_{0}=22$ C (see Supporting Material). For the reasonable value $K_{0}=0.7$ this results in $U_{0}\approx 0.52$ m/s, in fair agreement with our experimental findings. This is quite remarkable noting the number of approximations that enter the above derivation.
@@
To further support our findings, we have performed a series of simulations using an in-house boundary integral (BI) method, based on the code first described in [30]. The method has proven to be reliable for solving potential free surface flows and has been previously adapted to simulate drop impact [10]. For the numerical simulations, the flow inside the droplet is assumed to be a potential flow and the droplet shape to be axisymmetrical. This system is solved with the BI method to yield the droplet contour, where it is coupled to a compressible and condensable lubrication flow in the vapor layer. This is in contrast to the work in [10], where for the non-condensable air layer it had been coupled to an incompressible lubrication flow, and to [31], where a different phase change model was used. More details of the simulation technique and the numerical setup can be found in the Supporting Material.
We performed numerical simulations for similar conditions as the experimental ones reported in Fig. 2A, starting the simulation from an (arbitrary) height $h_{0}$, large enough to not influence results. More specifically, in Fig. 4A we extract the shape of the lower portion of an impacting droplet from the simulations. For each case, the first profile shown is at the moment that a sphere with the same speed and radius as the droplet would touch the surface in the absence of any gas or vapor layer. Subsequent profiles are shown in time intervals of $1.33\cdot 10^{-3}R_{0}/U_{0}$, such that, far from the substrate, successive profiles for different $U_{0}$ would be equidistant, and the last profile corresponds to when the simulation has stopped due to overlapping with the substrate.
Figure 4: (A) Time evolution of the axisymmetric and normalized profile $h/R_{0}$ as a function of the normalized radial coordinate $r/R_{0}$ for four boundary integral simulations of a HFE-7000 droplet impacting at different speeds $U_{0}=0.2$, $0.4$, $0.6$, $0.8$ m/s. The horizontal black dashed lines at $h=455$ nm indicate the height at which a droplet would enter the evanescent wave of the light source, and would thus show up in a TIR experiment. Color indicates time progression from yellow to dark blue. (B) Time evolution of the central dimple height $h_{d}$, normalized with the initial height $h_{0}$, for various impact velocities $U_{0}$ between $0.2$ m/s and $1.0$ m/s (blue to yellow). The straight dashed black line represents the time evolution of the position of the bottom of a moving sphere in the absence of vapor. Inset: Zoom into the blue dashed rectangular area. (C) The center touchdown time $\Delta t$ (blue pentagons), plotted as a function of the impact velocity $U_{0}$, where the continuous red curve is a hyperbolic fit to the data and the dashed red vertical line represents its asymptote. The red stars represent the entrapped vapor volume $V_{\text{e}}$ indicated in (A).
@@
At $U_{0}=0.6$ m/s, a qualitative difference appears in the dynamics of the bottom surface. From the profile sequence it becomes clear that also in this case the profile in the center is decelerated, but unlike what happened in the previous case, it continues to move downwards, and the bottom is nearly flat such that, even if some vapor is entrapped upon impact, it is likely to be very short lived: The subsequent build-up of pressure in the entrapped bubble will be able to quickly condense whatever vapor is left inside it. Finally, at the highest impact velocity ($U_{0}=0.8$ m/s in Fig. 4A) there is no sign of dimple formation at all: The droplet falls through the vapor layer as if no intermediate phase existed and hits the substrate with its center, without entrapment of a vapor bubble. Note that the spacing between profiles becomes larger with increasing speed, indicating considerably less deceleration for large $U_{0}$. All these findings are consistent with the experimental data presented in Fig. 2A. In the Supporting Material (Fig. S2) we compare the profiles of Fig. 4A to those of a non-condensable gas with the same properties.
@@
We now proceed to use the simulations to determine the transition impact velocity $U_{0,\text{thr}}$ from vapor entrapment to immediate touchdown of the liquid surface. To quantify $U_{0,\text{thr}}$, we determined the time difference $\Delta t$ between the moment of impact of the bottom of a sphere approaching the substrate in the absence of an intermediate phase and the impact time of the centre of the bottom surface of a simulated drop impact. To obtain the latter, we have tracked the dimple height $h_{d}$, which is the distance from the substrate to the bottom liquid surface at the center, i.e., at $r=0$, for simulations with impact velocities $0.2\leq U_{0}\leq 1$ m/s, as shown in Fig. 4B. As already mentioned, the simulations will stop a little before contact, such that we do not possess numerical data all the way to touchdown and linearly extrapolate the dimple height to determine $\Delta t$ as indicated in the inset of Fig. 4B.
@@
To complement the isothermal experiments, we have performed experiments with a significantly heated substrate, which in principle should counteract the condensation. The result is shown in Fig. 2B, for droplets impacting at $U_{0}\approx 0.7$ m/s. For superheat $\Delta T=T_{s}-T_{0}=2.4$ K, the entrapment is completely suppressed, as based on our isothermal experiments would be expected for this impact velocity. However, as soon as we increase the substrate temperature the droplet will start to entrap vapor again. For four degrees higher substrate temperature ($\Delta T=6.4$ K) the lifetime of the vapor bubble is extremely brief (less than 10 $\mu s$), the initial entrapment radius small, and moreover only visible in the first frame. At $\Delta T=11.3$ K, the entrapped vapor bubble remains visible in the first two frames and further incrementing $T_{s}$ to $\Delta T=17.6$ K we fully recover vapor entrapment: An axisymmetric wetted ring is found in the first recorded frame and the inner and outer contact lines expand as was previously observed in the isothermal low impact velocity experiments, where the entrapped vapor bubble remains visible up to the last frame shown in the figure.
@@
In summary, the process by which entrapment of a vapor bubble below a droplet impacting at higher velocities is suppressed, can be reversed by heating the substrate, frustrating the condensation that occurs at the vapor-liquid interface. The phase diagram in Fig. 5 traces the transition from vapor bubble entrapment to its suppression in the parameter space formed by the substrate superheat $\Delta T$ and the droplet impact velocity $U_{0}$. Black crosses denote impacts without entrapment and circles indicate that a vapor bubble was entrapped, where their color indicates the lifetime $t^{*}$ of the vapor bubble, changing from dark blue to bright yellow for increasing lifetime. For nearly isothermal impacts, no entrapment is recorded above $U_{0,\text{thr}}\approx 0.6$ m/s, in good agreement with the numerical simulations we conducted. When the superheat $\Delta T$ is increasing, this transition point keeps shifting to larger velocities. Notably, the entrapment lifetime along the transition line remains small, whereas points further away from the transition line live longer.
The superheat $\Delta T$ can be included by adding a parasitic linear heat conduction term through the vapor layer, $Sk_{V}(T_{s}-T_{V})/h\approx Sk_{V}\Delta T/h$ to the left hand side of (3). Using a second multiplicative numerical constant $Q_{0}$ in front of the parasitic heat conduction term, the condition $U_{cond}/U_{0}=1$ for $Q_{0}=0.9$ leads to the blue line in Fig. 5, which nicely demarcates the regions with and without vapor bubble entrapment in Fig. 5. Details of this calculation can be found in the Supporting Material.

### v2

@@ Abstract
In this work we replace gas by vapor, by focusing on the impact of boiling liquid droplets, that is, droplets that are at their boiling point, in thermal equilibrium with the vapor that surrounds them. More specifically, we want to focus on the entrapment of vapor inside the dimple formed below the droplet due to pressure accumulation. A related subject is that of droplets impacting in (or close to) the well-known Leidenfrost regime [24], where due to evaporation a droplet can levitate over a heated solid upon impact. These have been performed in atmospheric [30, 17, 8, 2, 7, 28] and reduced pressure conditions [31]. Unlike this previous work, we do not use large temperature differences to trigger rapid phase change, since the liquid will be in thermodynamical equilibrium with its own vapor, even before impact. When in such a state, minute energy exchanges of the droplet with its surroundings can trigger phase change [4]. Moreover, unlike the Leidenfrost impact work where phase change is limited to the evaporation of the liquid due to high substrate temperatures, in this boiling liquid state the condensation of vapor due to pressure locally rising above the vapor pressure of the liquid is also possible. We will show that the latter may lead to dramatic effects.
@@
In order to generate droplets in equilibrium with their own vapor and record their impact onto a substrate, we have developed a highly versatile setup, using HFE-7000 as a working fluid, which has an appreciable vapor pressure ($P_{V,0}=0.57$ bar) at room temperature. Here we provide a general description of the setup (see Fig. 1), while a more detailed description can be found in the Supporting Information (SI).
@@
First, a saturated liquid-vapor environment at temperature $T_{0}\approx 22$ C is created by allowing HFE-7000 to evaporate inside a hermetically sealed, vacated, and temperature-controlled chamber. Enough liquid is allowed into a reservoir (Res1) to ensure the chamber remains saturated with vapor during the entire experiment. Additional HFE-7000 liquid is allowed into a smaller reservoir (Res2), located higher in the setup, connected to a needle (N) with its tip slightly below the liquid level, allowing droplets to fall down and impact a sapphire prism (P). The height between needle and substrate can be varied to obtain an impact velocity $U_{0}$ between $0.2$ m/s and $1.5$ m/s. The substrate temperature $T_{s}$ is controlled between $T_{0}$ and $40$ C.
Side views of the impact are recorded at 16,000 fps and are used to measure the impact speed $U_{0}$ and initial radius $R_{0}$ of the droplets. Bottom views are captured at frame rates between $96,000$ and $200,000$ fps by illuminating the top surface of the prism internally with a collimated monochromatic light source (CLS2) from below, such that it totally reflects. If a droplet comes within the evanescent wave distance from the surface, then light is partly, and upon contact fully transmitted into the droplet, which results in dark regions on the otherwise bright surface. This technique is known as frustrated total internal reflection (TIR) and has been extensively used in the past [27, 14, 2].
@@
We have performed an extensive experimental campaign, divided in two parts. For the first part we have performed experiments where boiling droplets, falling at different velocities, impact a slightly heated surface. The latter is done to avoid the formation of a liquid film on top of the substrate that would interfere with the impact. For simplicity, we will refer to the experiments in this first part as (near-)isothermal experiments.
@@
Figure 2: (A) TIR image sequence (after background subtraction) of a series of HFE-7000 droplets with radius $R_{0}\approx 0.6$ $m$ in a vapor atmosphere at $T_{0}\approx 22$ C; the droplets impact a sapphire substrate heated to a temperature difference $\Delta T\lesssim 2$ C with impact velocities $U_{0}=0.24$, $0.38$, $0.51$, $0.74$ m/s, Fig. S2 and movies S1 to S5 in the SI. At low impact velocity a vapor bubble is entrapped, but at high impact velocities this entrapment is suppressed. (B) TIR image sequence (after background subtraction) of a series of droplets with radius $R_{0}\approx 0.6$ $m$ in a vapor atmosphere at $T_{0}\approx 22$ C impacting with velocity $U_{0}\approx 0.7$ m/s onto a sapphire substrate heated to an increasing temperature difference $\Delta T=2.4$, $6.4$, $11.3$, $17.6$ K (with respect to $T_{0}$), Movies S6 to S10 in the SI. At this impact velocity, the entrapment of vapor is suppressed for small $\Delta T$, but entrapment can be recovered by heating the substrate to a higher temperature. In all cases, time is measured with respect to the first TIR frame in which the droplet becomes visible.
@@
Figure 3: (A) Azimuthally averaged inner ($r_{b}$, blue points) and outer ($r_{d}$, orange points) radius evolution for an experiment at $U_{0}=0.38$ m/s, with images available in the SI, Fig. S1B. The orange line is obtained from a linear fit to $r_{d}^{2}$ plotted against $t$ and the blue line is a linear fit for $r_{b}$. The intersection of the two curves provides the initial radius $r_{0}$ and also the initial impact time. The entrapment time $t^{*}$ is then measured from the inner radius fit as the time it takes for $r_{b}(t)$ to go from $r_{0}$ to 0. (B) Initial entrapment radius $r_{0}$ as a function of the impact velocity $U_{0}$ for a set of experiments with a small temperature difference of the impact surface with respect to the ambient temperature ($\Delta T=1.77\pm 0.49$ K). The points represent the mean of a minimum of three repetitions and the bars display the total error. (C) entrapment time $t^{*}$ as a function of the impact velocity $U_{0}$, for the same set of experiments.
@@
In the next few paragraphs, through a simplified model, we find a quantitative explanation for the observed suppression of vapor bubble entrapment for increasing impact velocity. We begin considering an impacting droplet surrounded by a non-condensable gas. The pressure difference under such droplet is described by the lubrication flow of the gas as the droplet nears the surface [21, 1]. As argued in that work, we can estimate the pressure build up below the droplet from the axisymmetric Stokes equation in the lubrication layer as
@@
where $\mu_{g}$ is the dynamic viscosity of the gas, $h$ is the thickness and $\ell\sim\sqrt{R_{0}h}$ the radial extent of the lubrication layer, where the latter is estimated geometrically from the approach of a spherical droplet to a planar substrate, and the gas velocity $U_{g}$ is related to the impact velocity by continuity, $U_{g}h\sim U_{0}\ell$. By equating this $\Delta P$ with the pressure needed to decelerate the liquid, namely $\rho_{L}(\partial\phi/\partial t)\sim\rho_{L}U_{0}^{2}\ell/h$ (with $\phi$ the flow potential), one obtains a scaling relation for the dimple height $h_{d}$ in terms of the Stokes number $\mathit{St}\equiv\rho_{L}U_{0}R_{0}/\mu_{g}$
@@
For our case, however, the surrounding phase is vapor and this means that an increase in pressure can trigger condensation: As soon as the pressure in the vapor layer rises, the vapor will try to condense to return to the equilibrium state. For simplicity, let’s now consider a one-dimensional system composed of a solid substrate, an incoming liquid surface and an intermediate vapor layer. During pressure build up, vapor will try to condense and produce latent heat. Since simultaneously the temperature in the vapor rises, a thermal boundary layer will develop in the liquid, which will allow for transport of the latent heat into the liquid. Now, writing down the mass-energy balance at the liquid-vapor interface [23], an approximation we also used in [4, 5]
@@
where $R_{s}$ is the specific gas constant of the vapor phase and the equilibrium state variables of the vapor are denoted with subscript 0: $P_{V,0}$, $\rho_{V,0}$, and $T_{0}=T_{V,0}$. Finally, the boundary layer thickness $\delta_{th}$ can be approximated as $\delta_{th}\sim\sqrt{\pi\alpha_{L}\Delta t}$, with $\alpha_{L}$ the thermal diffusivity of the liquid and the time scale $\Delta t\approx\epsilon\frac{R_{0}}{U_{0}}$, where $\epsilon$ should be approximated from evaluating the time that the pressure rise in the lubrication layer starts being appreciable. This leads to $\epsilon\approx 0.02$, corresponding to $h\approx 0.02R_{0}$, which is justified in the SI and is both consistent with a similar estimate in [8] and with BI simulations (see next Section). Substituting everything into the expression for the condensation speed, and dividing by $U_{0}$ we obtain
@@
Now, the threshold condition $U_{cond}/U_{0}=1$ can be computed by turning the approximate result (8) into an equation using a multiplicative numerical constant $K_{0}$ in the right hand side and subsequently inserting literature values for the properties of HFE-7000 at $T_{0}=22$ C (see SI). For the reasonable value $K_{0}=1.6$ this results in $U_{0}\approx 0.52$ m/s, in fair agreement with our experimental findings. This is quite remarkable noting the number of approximations that enter the above derivation.
@@
To further support our findings, we have performed a series of simulations using an in-house boundary integral (BI) method, based on the code first described in [22]. The method has proven to be reliable for solving potential free surface flows and has been previously adapted to simulate drop impact [1]. For the numerical simulations, the flow inside the droplet is assumed to be a potential flow and the droplet shape to be axisymmetrical. This system is solved with the BI method to yield the time evolution of the droplet contour, where it is coupled to a compressible and condensable lubrication flow in the vapor layer. This is in contrast to the work in [1], where for the non-condensable air layer it had been coupled to an incompressible lubrication flow, and to [10], where a different phase change model was used. More details of the simulation technique and the numerical setup can be found in the SI.
We performed numerical simulations for similar conditions as the experimental ones reported in Fig. 2A, starting the simulation from an (arbitrary) height $h_{0}$, large enough to not influence results. More specifically, in Fig. 4A we extract the shape of the lower portion of an impacting droplet from the simulations. For each case, subsequent profiles are shown in time intervals of $1.33\cdot 10^{-3}R_{0}/U_{0}$, such that, far from the substrate, successive profiles for different $U_{0}$ would be equidistant, and the last profile corresponds to just before the simulation has stopped due to overlapping with the substrate.
Figure 4: (A) Time evolution of the axisymmetric and normalized profile $h/R_{0}$ as a function of the normalized radial coordinate $r/R_{0}$ for four boundary integral simulations of a HFE-7000 droplet impacting at different speeds $U_{0}=0.2$, $0.4$, $0.6$, $0.8$ m/s. The horizontal black dashed lines at $h=455$ nm indicate the height at which a droplet would enter the evanescent wave of the light source, and would thus show up in a TIR experiment. Color indicates time progression from yellow to dark blue. (B) Time evolution of the center height $h_{c}$, normalized with the initial height $h_{0}$, for various impact velocities $U_{0}$ between $0.2$ m/s and $1.0$ m/s (blue to yellow). The straight dashed black line represents the time evolution of the position of the bottom of a moving sphere in the absence of vapor. Inset: Zoom into the blue dashed rectangular area. (C) The center touchdown time $\Delta t$ (blue pentagons), plotted as a function of the impact velocity $U_{0}$, where the continuous red curve is a hyperbolic fit to the data and the dashed red vertical line represents its asymptote. The red stars represent the entrapped vapor volume $V_{\text{e}}$ indicated in (A).
@@
At $U_{0}=0.6$ m/s, a qualitative difference appears in the dynamics of the bottom surface. From the profile sequence it becomes clear that also in this case the profile in the center is decelerated, but unlike what happened in the previous case, it continues to move downwards, and the bottom is nearly flat such that, even if some vapor is entrapped upon impact, it is likely to be very short lived: The subsequent build-up of pressure in the entrapped bubble will be able to quickly condense whatever vapor is left inside it. Finally, at the highest impact velocity ($U_{0}=0.8$ m/s in Fig. 4A) there is no sign of dimple formation at all: The droplet falls through the vapor layer as if no intermediate phase existed and hits the substrate with its center, without entrapment of a vapor bubble. Note that the spacing between profiles becomes larger with increasing speed, indicating considerably less deceleration for large $U_{0}$. All these findings are consistent with the experimental data presented in Fig. 2A. In the SI (Fig. S3) we compare the profiles of Fig. 4A to those of a non-condensable gas with the same properties.
@@
We now proceed to use the simulations to determine the transition impact velocity $U_{0,\text{thr}}$ from vapor entrapment to immediate touchdown of the liquid surface. To quantify $U_{0,\text{thr}}$, we determined the time difference $\Delta t$ between the moment of impact of the bottom of a sphere approaching the substrate in the absence of an intermediate phase and the impact time of the centre of the bottom surface of a simulated drop impact. To obtain the latter, we have tracked the center height $h_{c}$, which is the distance from the substrate to the bottom liquid surface at the center, i.e., at $r=0$, for simulations with impact velocities $0.2\leq U_{0}\leq 1$ m/s, as shown in Fig. 4B. As already mentioned, the simulations will stop a little before contact, such that we do not possess numerical data all the way to touchdown and linearly extrapolate the dimple height to determine $\Delta t$ as indicated in the inset of Fig. 4B.
@@
To complement the isothermal experiments, we have performed experiments with a significantly heated substrate, which in principle should counteract the condensation. This is reminiscent of the well-known Leidenfrost effect [24], where a droplet is made to hover over a strongly heated substrate levitated by its own vapor. The result is shown in Fig. 2B, for droplets impacting at $U_{0}\approx 0.7$ m/s. For superheat $\Delta T=T_{s}-T_{0}=2.4$ K, the entrapment is completely suppressed, as based on our isothermal experiments would be expected for this impact velocity. However, as soon as we increase the substrate temperature the droplet will start to entrap vapor again. For four degrees higher substrate temperature ($\Delta T=6.4$ K) the lifetime of the vapor bubble is extremely brief (less than 10 $\mu s$), the initial entrapment radius small, and moreover only visible in the first frame. At $\Delta T=11.3$ K, the entrapped vapor bubble remains visible in the first two frames and further incrementing $T_{s}$ to $\Delta T=17.6$ K we fully recover vapor entrapment: An axisymmetric wetted ring is found in the first recorded frame and the inner and outer contact lines expand as was previously observed in the isothermal low impact velocity experiments, where the entrapped vapor bubble remains visible up to the last frame shown in the figure.
@@
In summary, the process by which entrapment of a vapor bubble below a droplet impacting at higher velocities is suppressed, can be reversed by heating the substrate, frustrating the condensation that occurs at the vapor-liquid interface. The phase diagram in Fig. 5 traces the transition from vapor bubble entrapment to its suppression in the parameter space formed by the substrate superheat $\Delta T$ and the droplet impact velocity $U_{0}$. Black crosses denote impacts without entrapment and circles indicate that a vapor bubble was entrapped, where their color represents the lifetime $t^{*}$ of the vapor bubble, changing from dark blue to bright yellow for increasing lifetime. For nearly isothermal impacts, no entrapment is recorded above $U_{0,\text{thr}}\approx 0.6$ m/s, in good agreement with the numerical simulations we conducted. When the superheat $\Delta T$ is increasing, this transition point keeps shifting to larger velocities. Notably, the entrapment lifetime along the transition line remains small, whereas bubbles corresponding to points further away from the transition line live longer.
Inspired by the analysis of the ordinary [24] and dynamic [8] Leidenfrost phenomenon, the superheat $\Delta T$ can be included by adding a parasitic linear heat conduction term through the vapor layer, $Sk_{V}(T_{s}-T_{V})/h\approx Sk_{V}\Delta T/h$ to the left hand side of (3), where $h=h_{d}$ is again estimated from the dimple condition (2). Using a second multiplicative numerical constant $Q_{0}$ in front of the parasitic heat conduction term, the condition $U_{cond}/U_{0}=1$ for $Q_{0}=1.35$ leads to the blue line in Fig. 5, which nicely demarcates the regions with and without vapor bubble entrapment in Fig. 5. Details of this calculation can be found in the SI.